Qd=1200-2P and Qs-3p what is the equilibrium price and quantitu


1
Expert's answer
2022-02-15T11:06:59-0500

Equilibrium price("E_{p}")

"E_{p}=Q_{s}=Q_{d}"

"Q_{s}=Q_{d}"

"3_{P}=1200-2_{P}"

"3_{P}+2_{P}=1200"

"5_{P}=1200"

"\frac{5_{P}}{5}=\frac{1200}{5}"

"p=240"


Quantity("Q")

"Q=3_{P}"

"Q=3\times240"

"Q=720"